# Metastability induced by non-reciprocal adaptive couplings in Kuramoto models

### v1

@@ Abstract
Non-reciprocal couplings are frequently found in systems out-of-equilibrium such as neuronal networks. We consider generalized Kuramoto models with non-reciprocal adaptive couplings. The non-reciprocity refers to the type of couplings according to Hebbian or anti-Hebbian rules and to different time scales, on which the couplings evolve. The main effect of this specific combination of deterministic dynamics is an induced metastability of anti-phase synchronized clusters of oscillators. Metastable switching is typical for neuronal networks and a characteristic of brain dynamics. We analyze the metatstability as a function of the system parameters, in particular of the size and the network connectivity. The mechanism behind sudden changes in the order parameters are individual oscillators which change their cluster affiliation from time to time, providing “weak ties” between clusters of synchronized oscillators. The time series have random features, but derive from deterministic dynamics.
@@ I Introduction
Among complex networks, adaptive networks have attracted much attention over the last years. Adaptation and learning are essential processes in evolution. In chemical networks, the reaction rates may dynamically be adapted to improve the productivity. In social systems, agents may adapt their strategies to those of their friends or imitate the strategies of successful agents. In technical applications, adaptive couplings are used to prevent system´s failures. In relation to neuronal networks, one of the first studies of plasticity and learning in a network of coupled phase oscillators is the work of [1] and somewhat later the work of [2]. More recently, the role of plasticity and adaptation is discussed in a review of [3]. Modular structure on meso- and macroscales can emerge when adaptively reinforced synchronization competes with a constraint on nodes to establish connections with other units of the network [4]. Adaptation is also considered for adaptive multiplex networks, see, for example, the work of [5, 6, 7]. Adaptive coupling can have the interesting effect of inducing collective excitability and self-sustained bursting oscillations in globally coupled populations of non-excitable units [8]. Insights from neural networks with adaptive couplings can be exploited to explain dynamical properties of multi-frequency clusters in power grid networks, described by Kuramoto-Sakaguchi phase oscillators with inertia [9]. The fundamental relation between power grid and neural networks is established in [10]; according to this relation phase oscillator models with inertia correspond to a particular class of adaptive networks.
In this paper we will consider special cases of the work of [11, 12, 13]. According to the work of [11], a multilayer structure and chimera states emerge in a self-organized way if generalized globally coupled Kuramoto models are considered with adaptive coupling. Subnetworks of densely coupled elements form sequentially whose size is hierarchically ordered and which decouple as result of the hierarchical structure. If the global coupling in such networks is replaced by a random sparse structure of connections, chimera states form which retain features of a hierarchical organization, and the set of elements that form coherent groups can be rearranged during the network evolution [12]. When the couplings adapt and have their own dynamics, it amounts to plasticity in relation to neuronal networks. In particular, the couplings may change as a function of the timing between the oscillators. If spike-time dependent plasticity enters also the natural frequencies of the oscillators, heterogeneous layered clusters with different frequencies show up from homogeneous oscillator populations [13]. For oscillatory neuronal populations with spike-timing dependent plasticity it is shown in [14] that for an optimal noise level the amount of synaptic coupling gets maximal, leading to noise-induced self-organized synaptic connectivity. Here, the application of noise counteracts desynchronization that would naively be expected as result of noise.
At another frontier, non-reciprocity is topical in systems out-of-equilibrium, since the validity of Newton‘s third law (for every action is an equal and opposite reaction) is the exception rather than the rule in these systems. Arguably the most plausible example is from social systems where a relation of friendship holds in one direction but more enmity than friendship in the opposite direction. Populations with conformist and contrarian members are considered in [15]. Non-reciprocity is also found in active matter [16, 17, 18], metamaterials [19, 20], or in relation to game theory [21]. Occasionally it is discussed in neuronal networks [1, 22]. In [1] asymmetric networks are considered, where only part of the synaptic connections between pairs of neurons are non-reciprocal and have slow dynamic response with the effect that temporal association becomes possible as recall of time sequences and cycles of patterns. The work of [22] introduces two populations of excitatory and inhibitory populations in a Kuramoto model with the non-reciprocal feature that each oscillator of the excitatory (inhibitory) population exerts a positive (negative) influence on each oscillator of the inhibitory (excitatory) population, respectively. This way, the Kuramoto model accounts for the onset of excitatory/inhibitory-based neuronal rhythms.
Non-reciprocity can have strong implications on the collective behavior of many-body systems. It may lead to time-dependent phases in which spontaneously broken continuous symmetries are dynamically restored [23]. In view of general manifestations of non-reciprocity, we consider systems with self-organizing structures, of which synchronization is a prototypical form. More specifically, this paper is on non-reciprocity in adaptive networks and its impact on synchronization patterns. The dynamics of nonreciprocal adaptation may be realized in many ways, via the type of favored interaction (alignment or anti-alignment, Hebbian or anti-Hebbian rules), via the choice of time scales, via feedback as in [1] or frequency adaptation to name some of them.
In this paper we focus on generalized Kuramoto models with both non-reciprocal and adaptive couplings. When we combine Hebbian and anti-Hebbian adaptation rules for the couplings, acting on quite different time scales, we observe switching dynamics between metastable states, characterized by anti-phase cluster synchronization (exceptionally also full synchronization) unless the choice of parameters leads to incoherent oscillations.
Many experimental hints exist that typical states in the brain are metastable [24]. The structure of these metastable states is reflected in functional neuroimage experiments. In particular cognitive phenomena rely on transient dynamics such as working memory and decision making [25, 26]. The idea is that brain activity is organized in spatiotemporal patterns through transient metastable states [27, 28]. For a long time the search was for a dynamic framework that describes transient reproducible and metastable dynamics. One such framework is provided by heteroclinic dynamics, for a recent review see [29]. For a review on metastable dynamics see, for example, [30].
In this paper, we demonstrate how -differently from heteroclinic dynamics- a set of deterministic equations can lead to switching dynamics, here induced by a fast change between attractive and repulsive couplings in half of the node connections, and a slow change between both types in the other half. In detail we consider an all-to-all coupling topology, but we present also results for a random network, for which we vary the density of connectivity. We briefly consider effects of non-reciprocal adaptivity when also the natural frequencies become coupling dependent.
The paper is organized as follows: In section 2 we introduce different versions of the model. Section 3 presents the results in terms of synchronization patterns as function of the model parameters, in particular of the time scales, the system size and the grid topology. Section 4 contains our conclusions. In the appendix we provide an analytical understanding of the increasing instability with the system size.
@@
Figure 1: Categorical heatmap of collective states as a function of $(\varepsilon_{1},\varepsilon_{2})$ for (a) $N=10$ oscillators and (b) $N=20$ oscillators, averaged over 55 independent realizations. Colors indicate two anti-phase clusters (orange), and incoherent (gray) states; $\epsilon_{1}$ is varying within [0, 0.1] and $\epsilon_{2}$ is varying within [0, 0.003] with a step length of 0.0001. The red dashed curve should guide the eyes and roughly marks the transition boundary between two anti-phase clusters and incoherent states via a quadratic least square fit.
@@ II The Model
where $\omega_{i}$ denotes the intrinsic natural frequency, and $k_{ij}$ is the time-dependent coupling strength. The term $\xi_{i}(t)$ represents additive Gaussian white noise of intensity $\sigma$:
@@
The coupling matrix $\mathbf{k}=[k_{ij}]$ evolves as
We choose $\beta_{1}=-\frac{\pi}{2}$ and $\beta_{2}=+\frac{\pi}{2}$, and $\varepsilon_{1},\varepsilon_{2}>0$ are the adaptation rates for the upper and lower triangular parts of the matrix. The diagonal elements of the coupling matrix, $k_{ii}$, are set to zero to avoid self-coupling. This means that for connections from a lower-index oscillator $i$ to a higher-index oscillator $j$, the coupling tends to adjust with a phase lag of $\pi/2$ (Hebbian plasticity rule). For connections in the opposite direction, the adjustment follows a phase lead of $\pi/2$ (anti-Hebbian rule). For simplicity the upper-triangular (lower-triangular) couplings $k_{ij}$ with $i<j$ ($i>j$) are termed the upper (lower) couplings in the following. Throughout the paper we will keep the slow (fast) time scale associated with the anti-Hebbian (Hebbian) rule, respectively, since otherwise we observe only incoherent oscillations.
In most previous studies [31]-[32] the adaptation rules were chosen symmetric with $\varepsilon_{1}=\varepsilon_{2}>0$. In these cases, if the phase shift was $\beta_{1}=\beta_{2}=+\pi/2$, the system typically remained incoherent [11], and if the phase shift was $\beta_{1}=\beta_{2}=-\pi/2$ , two anti-phase clusters could form [11].
As stochastic extension of the standard Euler method, Equations (1) and (3) are integrated numerically using the Euler–Maruyama scheme with a fixed time step $\Delta t=0.01$ and $\theta_{i}\mapsto\theta_{i}\bmod 2\pi$.
The initial phases $\theta_{i}(0)$ are sampled uniformly from $0,2\pi)$, while (unless stated otherwise) the initial couplings $k_{ij}(0)$ are drawn from a uniform distribution in [1,1][-1,1] with $k_{ii}(0)=0$. The natural frequencies $\omega_{i}$ are in most cases set to zero, otherwise sampled from a normal distribution with a fixed mean $\omega_{\mathrm{mean}}$ and a fixed standard deviation $\omega_{\mathrm{std}}$.
The degree of phase coherence is quantified using two complex quantities, the first and second Kuramoto order parameters:
@@
where $R\in[0,1]$ measures the global synchronization level, and $R_{2}\in[0,1]$ captures the presence of two anti-phase clusters. The limiting cases are as follows: $R\approx 1$ and $R_{2}\approx 1$ for complete in-phase synchronization; $R\approx 0$ and $R_{2}\approx 1$ for two anti-phase clusters; and both $R$ and $R_{2}$ small for incoherent dynamics.
After an initial transient, the long-time averaged order parameters $\langle R\rangle$ and $\langle R_{2}\rangle$ are computed over a sufficiently large portion of the simulation succeeding the transient. As we will see, for our results it is sufficient to classify the collective states as follows: if $\langle R\rangle$ and $\langle R_{2}\rangle$ both are greater than a threshold $R_{\mathrm{th}}$, the system is in the in-phase synchronized state. If $\langle R\rangle$ is less than or equal to $R_{\mathrm{th}}$, and $\langle R_{2}\rangle$ is greater than $R_{\mathrm{th}}$, the system is in the two anti-phase clusters state. In all other cases, the system is considered incoherent. For the threshold we have chosen $R_{\mathrm{th}}=0.8$, though usually for anti-phase synchronized clusters $R_{2}>0.8$.

### v2

@@ Abstract
Non-reciprocal couplings are frequently found in systems out-of-equilibrium such as neuronal networks. We consider generalized Kuramoto models with non-reciprocal adaptive couplings. The non-reciprocity refers to the type of couplings according to Hebbian or anti-Hebbian rules and to different time scales on which the couplings evolve. The main effect of this specific combination of deterministic dynamics is an induced metastability of anti-phase synchronized clusters of oscillators. Metastable switching is typical for neuronal networks and a characteristic of brain dynamics. We analyze the metastability as a function of the system parameters, in particular of the size and the network connectivity. The mechanism behind sudden changes in the order parameters is individual oscillators which change their cluster affiliation from time to time, providing “weak ties” between clusters of synchronized oscillators, where an individual oscillator may represent an entire brain area. The time series exhibit random features but arise from deterministic dynamics.
@@ I Introduction
To bridge the gap between behavioral and neural views of learning, neural analogues of behavioral modification paradigms have been postulated. An example of such an attempt was D. Hebb’s suggestion in 1949 [1] that when a cell $A$ repeatedly and persistently takes part in firing another cell $B$, $A^{\prime}s$ efficiency in firing $B$ is increased. This is a rule for synaptic plasticity. It amounts to an analogue of associative conditioning. It is then typical for neural network models to employ a presynaptic and a postsynaptic correlation rule for altering connectivity as a mathematical representation of Hebb’s postulate. Inspired by Hebb’s seminal work, Rosenblatt proposed the perceptron [2] to establish a relation between biophysics and psychology. His goal was to predict learning curves from neurological variables and vice versa, neurological variables from learning curves. He proposed a quantitative statistical approach to the organization of cognitive systems. This was in 1958. Already in the seventies and eighties, many adaptive neural network theories were considered in which neuron-like adaptive elements behave as single unit analogues of associative conditioning. In particular, such an adaptive element was proposed in [3] that respects facts of animal learning theory. It implements an essential feature of classical conditioning that the element learns to increase its response rate in anticipation of increased stimulation. Later many theoretical studies of learning in neural systems were based on the Hopfield model [4] and McCulloch-Pitts units [5] or “integrate-and-fire” models of neurons, coupled via synapses that accumulate the presynaptic activity into an increasing membrane potential of neurons which fire when a certain threshold is crossed [6]. The synaptic strength there is not fixed.
In a complementary approach, neuronal activity is modeled by periodic oscillators which replace the neurons or entire sub-populations of neurons such as cortical areas. The reason is that rhythmic activity is observed, for example, in central pattern generators, visual or olfactory systems [7]. When neural networks are described as networks of coupled oscillators, the role of relative spike timing is played by the individual oscillator phases. After all, one can reduce the oscillator dynamics to a pure phase dynamics if the coupling is assumed to be weak. This led to the proposal of Kuramoto’s model of phase oscillators [8], for a later review see [9]. In its original version, globally coupled Kuramoto oscillators with nonidentical frequencies exhibit a second-order phase transition to a synchronized state as the coupling strength increases above a critical value. Various generalizations of the Kuramoto model have been considered afterwards.
@@
One of the first studies of plasticity and learning in a network of coupled phase oscillators is the work of [10] and somewhat later the work of [11]. Coupling adaptation in oscillatory systems leads to many interesting effects also beyond applications to neural networks, since the level of description is abstract enough to allow different interpretations. Insights from neural networks with adaptive couplings can be exploited to explain dynamical properties of multi-frequency clusters in power grid networks, described by Kuramoto-Sakaguchi phase oscillators with inertia [12]. The fundamental relation between power grid and neural networks is established in [13]; according to this relation phase oscillator models with inertia correspond to a particular class of adaptive networks. Modular structure on meso- and macroscales can emerge when adaptively reinforced synchronization competes with a constraint on nodes to establish connections with other units of the network [14]. Adaptation is also considered for adaptive multiplex networks, see, for example, the work of [15, 16, 17]. Adaptive coupling can have the interesting effect of inducing collective excitability and self-sustained bursting oscillations in globally coupled populations of non-excitable units [18]. For oscillatory neuronal populations with spike-timing dependent plasticity it is shown in [19] that for an optimal noise level the amount of synaptic coupling gets maximal, leading to noise-induced self-organized synaptic connectivity. Here, the application of noise counteracts desynchronization that would naively be expected as result of noise.
@@
Figure 1: Categorical heatmap of collective states as a function of $(\varepsilon_{1},\varepsilon_{2})$ for (a) $N=10$ oscillators and (b) $N=20$ oscillators, averaged over $200$ independent realizations. For a specific choice of $\varepsilon_{1},\varepsilon_{2}$ colors indicate the state to which the system converges: orange for two anti-phase clusters, and gray for incoherent states; $\varepsilon_{1}$ is varying within [0, 0.01] and $\varepsilon_{2}$ is varying within [0, 0.003] with a step length of 0.0001. The red dashed curve serves to guide the eyes and roughly marks the transition boundary between two anti-phase clusters and incoherent states via a quadratic least square fit.
According to the work of [20], a multilayer structure and chimera states emerge in a self-organized way if generalized globally coupled Kuramoto models are considered with adaptive couplings. Subnetworks of densely coupled elements form sequentially whose size is hierarchically ordered and which decouple as result of the hierarchical structure. If the global coupling in such networks is replaced by a random sparse structure of connections, chimera states form which retain features of a hierarchical organization, and the set of elements that form coherent groups can be rearranged during the network evolution [21]. When the couplings adapt and have their own dynamics, it amounts to plasticity in relation to neuronal networks. In particular, the couplings may change as a function of the timing between the oscillators. If spike-time dependent plasticity also enters into the natural frequencies of the oscillators, heterogeneous layered clusters with different frequencies show up from homogeneous oscillator populations [22]. In this paper we will consider special cases of the work of [20, 21, 22].
More recently, possible biological or neurological mechanisms related to learning, plasticity, and adaptation are reviewed in [23]. Hebbian learning is one important mechanism of brain plasticity, but Hebbian learning alone would lead to dynamic instability and requires compensatory processes on multiple timescales [24]. Various mechanisms are listed in [23] to achieve dynamic stability. It seems still to be relatively unknown at which spatial and temporal scales Hebbian, homeostatic, and other plasticity mechanisms interact, and which functional role they play. In this work, we will consider two plasticity mechanisms, Hebbian and anti-Hebbian, acting at quite different time scales, but this choice is not motivated by neurological experiments, it serves to explore the effect of non-reciprocity.
At another and very different frontier, non-reciprocity is topical in systems out-of-equilibrium, since the validity of Newton‘s third law (for every action is an equal and opposite reaction) is the exception rather than the rule in these systems. Arguably the most plausible example is from social systems where a relation of friendship holds in one direction but more enmity than friendship in the opposite direction. Populations with conformist and contrarian members are considered in [25]. Non-reciprocity is also found in active matter [26, 27, 28], metamaterials [29, 30], or in relation to game theory [31]. Occasionally it is discussed in neuronal networks [10, 32]. In [10], asymmetric networks are considered, where only part of the synaptic connections between pairs of neurons are non-reciprocal and have slow dynamic response with the effect that temporal association becomes possible as recall of time sequences and cycles of patterns. The work of [32] introduces two populations of excitatory and inhibitory populations in a Kuramoto model with the non-reciprocal feature that each oscillator of the excitatory (inhibitory) population exerts a positive (negative) influence on each oscillator of the inhibitory (excitatory) population, respectively. This way, the Kuramoto model accounts for the onset of excitatory/inhibitory-based neuronal rhythms.
Particularly in view of collective behavior, non-reciprocity can have strong implications on many-body systems. It may lead to time-dependent phases in which spontaneously broken continuous symmetries are dynamically restored [33]. In view of general manifestations of non-reciprocity, we consider systems with self-organizing structures, of which synchronization is a prototypical form. More specifically, this paper is on non-reciprocity in adaptive networks and its impact on synchronization patterns. The dynamics of non-reciprocal adaptation may be realized in many ways, via the type of favored interaction (alignment or anti-alignment, Hebbian or anti-Hebbian rules), via the choice of time scales, via feedback as in [10] or frequency adaptation to name some of them.
In this paper, we focus on generalized Kuramoto models with both adaptive and non-reciprocal couplings. When we combine Hebbian and anti-Hebbian adaptation rules for the couplings, acting on quite different time scales, we observe switching dynamics between metastable states, characterized by anti-phase cluster synchronization (exceptionally also full synchronization) unless the choice of parameters leads to incoherent oscillations. Emerging metastability is the main novel feature of our results. Therefore, the results may be relevant to neural network applications, since many experimental hints exist that typical states in the brain are metastable [34]. The structure of these metastable states is reflected in functional neuroimaging experiments. In particular cognitive phenomena rely on transient dynamics such as working memory and decision making [35, 36]. The idea is that brain activity is organized in spatiotemporal patterns through transient metastable states [37, 38]. For a long time the search was for a dynamic framework that describes transient reproducible and metastable dynamics. One such framework is provided by heteroclinic dynamics, for a recent review see [39]. For a review on metastable dynamics see, for example, [40].
In this paper, we demonstrate how -differently from heteroclinic dynamics- a set of deterministic equations can also lead to switching dynamics, here induced by a fast change between attractive and repulsive couplings in half of the node connections, and a slow change between both types in the other half. In detail we consider an all-to-all coupling topology, but we also present results for a random network, for which we vary the density of connectivity. We briefly consider effects of non-reciprocal adaptivity when also the natural frequencies become coupling dependent.
The paper is organized as follows: In section 2 we introduce different versions of the model. Section 3 presents the results in terms of synchronization patterns as function of the model parameters, in particular of the time scales, the system size, and the grid topology. Section 4 contains our conclusions. In the appendix we provide an analytical understanding of the increasing instability with the system size.
@@ II The Model
where $\omega_{i}$ denotes the intrinsic natural frequency, and $k_{ij}$ is the time-dependent coupling strength, with i,j=1,,Ni,j=1,...,N. The coupling is all-to-all and directed, and the matrix elements $\mathbf{k}=[k_{ij}]$ evolve as
We choose $\beta_{1}=-\frac{\pi}{2}$ and $\beta_{2}=+\frac{\pi}{2}$. The parameters $\varepsilon_{1},\varepsilon_{2}>0$ set the different time scales for the adaptation rates for the upper and lower triangular parts of the matrix. The diagonal elements of the coupling matrix, $k_{ii}$, are set to zero to avoid self-coupling. This means that for connections from a lower-index oscillator $i$ to a higher-index oscillator $j$, the coupling tends to adjust with a phase lag of $\pi/2$ (Hebbian plasticity rule, where the weights for a pair of in-phase (or anti-phase) oscillators will increase (or decrease), respectively,) see, for example, [22]. For connections in the opposite direction, the adjustment follows a phase lead of $\pi/2$ (anti-Hebbian rule), discussed, for example, in [20]. For simplicity the upper-triangular (lower-triangular) couplings $k_{ij}$ with $i<j$ ($i>j$) are termed the upper (lower) couplings in the following. The term $\xi_{i}(t)$ represents additive Gaussian white noise of intensity $\sigma$:
@@
Throughout the paper we will keep the slow (fast) time scale associated with the anti-Hebbian (Hebbian) rule, respectively, since otherwise we observe only incoherent oscillations. Our motivation for this choice does not directly come from experiments, it is to test the effect of non-reciprocity when it is implemented in the adaptation dynamics by the type and the time scales of adaptation. Moreover, for $\epsilon_{1}=\epsilon_{2}$ it allows a comparison to the results of the work of [41, 20, 21, 22] for special choices of their parameters. For example, when the phase dynamics includes an additional parameter $\alpha$ to simulate a phase lag of the interaction, we set it to zero unless otherwise stated. In previous studies of [42]-[43], the adaptation rules were chosen symmetric with $\varepsilon_{1}=\varepsilon_{2}>0$. In these cases, if the phase shift was $\beta_{1}=\beta_{2}=+\pi/2$, the system typically remained incoherent [20], while if the phase shift was $\beta_{1}=\beta_{2}=-\pi/2$ , two anti-phase clusters could form [20].
As a stochastic extension of the standard Euler method, Eqs. (1) and (2) are integrated numerically using the Euler–Maruyama scheme with a fixed time step $\Delta t=0.01$ and $\theta_{i}\mapsto\theta_{i}\bmod 2\pi$.
The initial phases $\theta_{i}(0)$ are sampled uniformly from $0,2\pi)$, while (unless stated otherwise) the initial couplings $k_{ij}(0)$ are drawn from a uniform distribution in [1,1][-1,1] with $k_{ii}(0)=0$. The natural frequencies $\omega_{i}$ are in most cases set to zero, for some test runs (not presented) we sampled them from a normal distribution with a fixed mean $\omega_{\mathrm{mean}}$ and a fixed standard deviation $\omega_{\mathrm{std}}$.
The degree of phase coherence is quantified using two complex quantities, the first Kuramoto order parameter $R$ [8, 9, 44] and the second Kuramoto order parameter $R_{2}$ [44]:
@@
where $R\in[0,1]$ measures the global synchronization level, and $R_{2}\in[0,1]$ captures the presence of two anti-phase clusters [41]. It is particularly useful to characterize our synchronization patterns. The limiting cases are as follows: $R\approx 1$ and $R_{2}\approx 1$ for complete in-phase synchronization; $R\approx 0$ and $R_{2}\approx 1$ for two anti-phase clusters; and both $R$ and $R_{2}$ small for incoherent dynamics.
Figure 3: Snapshots of the system state before and after a switching event for $N=20$. (a–c) State at $t=2000$ (before the switch). (a) Instantaneous phases on the unit circle showing two well-separated clusters. (b) Time evolution of the order parameters $R$ and $R_{2}$. (c) Time evolution of the cluster sizes, showing a partition into $14$ and 66 oscillators. (d–f) State at $t=4000$ (after the switch). (d) Instantaneous phases. (e) Evolution of order parameters; note the jump in $R$ and $R_{2}$ coinciding with the switch. (f) Cluster sizes reorganize into a partition of $16$ and 44 oscillators.
After an initial transient, the long-time averaged order parameters $\langle R\rangle$ and $\langle R_{2}\rangle$ are computed over a sufficiently large portion of the simulation succeeding the transient. Unless otherwise specified, we simulate the system for a total of $T_{total}=1.8\times 10^{6}$ time steps. To eliminate initial transient behaviors, we discard the first $30\%$ of the simulation (approx. $5.4\times 10^{5}$ steps) before computing the long-time averaged order parameters $\langle R\rangle$ and $\langle R_{2}\rangle$. As we will see, for our results it is sufficient to classify the collective states as follows: if $\langle R\rangle$ and $\langle R_{2}\rangle$ both are greater than a threshold $R_{\mathrm{th}}$, the system is in the in-phase synchronized state. If $\langle R\rangle$ is less than or equal to $R_{\mathrm{th}}$, and $\langle R_{2}\rangle$ is greater than $R_{\mathrm{th}}$, the system is in the two anti-phase clusters state. In all other cases, the system is considered incoherent. For the threshold we have chosen $R_{\mathrm{th}}=0.8$, though usually for anti-phase synchronized clusters $R_{2}>0.8$. The threshold of $R_{th}=0.8$ acts as a conservative lower bound [45, 46]. In the metastable states that we observe, the synchronized clusters are typically tightly locked, yielding $R_{2}$ values well above $0.9$. A threshold of $0.8$ reliably separates these coherent states from incoherent regimes while allowing for small fluctuations induced by the adaptive, non-reciprocal coupling dynamics.